OpenStudy (hesan):

two cards are drawn at random from a deck of well shuffled cards. Determine the probability that both the cards are Aces.

OpenStudy (anonymous):

this is what i got as the answer: (4/52) * (3/51) = 1/(13*17) = 1/ 221

OpenStudy (ganpat):

4c2 / 52c 2 = (4 * 3 * 2 )/ (2 * 2) / (52 * 51 * 50!) / (2 * 50 !) = 6 / (26 * 51) = 1/ 221..